Grade 9Mechanics

The heart is pumping blood at x kg per unit time, with constant velocity v. The force needed

Mohammed Areeb
8 Years agoGrade 9

2 Answers

Sanju
8 Years ago
Force = Rate of change of momentum 
F = mv/t = (m/t)v = xv      [Note: m/t is x, which is given in the question]
So the required force is xv
elon musk
6 Years ago
newtons first law states that “ if a body is at constant velocity the net FORCE is zero”. so if the heart pumps at constant velocity the force must be equal to zero.